# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v2

@@ I.1 Main Results
We find that the single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, up to allowed error $\varepsilon\in(0,1)$ under the action of Gibbs preserving superchannels, are $\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$, respectively (Theorem 2). We have $F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$, which provide exact operational meaning to the resource-theoretic hypothesis-testing free energy and the max-free energy of a quantum channel. The asymptotic athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, under parallel uses of channels, are equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ (Theorem 3). The resource-theoretic free energy of a channel $\mathcal{N}$, $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, and thus its asymptotic athermality distillation and formation rates are both equal to $\frac{\beta}{2}F^{\beta}[\mathcal{N}]$, providing the exact operational interpretation to the resource-theoretic free energy of a quantum channel. Both asymptotic athermality distillation and formation rates of a quantum channel are the same, implying that the athermality of quantum channels under Gibbs preserving superchannels is (asymptotically) reversible.
@@ II Preliminaries
The generalized channel divergences are monotonically nonincreasing under the action of quantum superchannels [30, Theorem 1]. If the generalized state divergence is nonnegative, then the corresponding channel divergence is also nonnegative. If the generalized state divergence is faithful, i.e., minimum if and only the states are equal, then the corresponding channel divergence is also faithful, i.e., attains minimum if and only if the channels are equal.
@@ III.1 Properties of the dynamical free energy
All the generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ satisfy (A1) and (A2) readily, whenever $\beta^{-1}\mathbf{D}(\rho\|\gamma^{\beta})$ represents some resource-theoretic free energy function $f(\rho)$ for states $\rho$ [12, 23]. The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ remains invariant under the action of Gibbs preserving unitary superchannels. Faithfulness (A3) is satisfied by the generalized free energy whose generalized channel divergence is faithful. For all $\alpha\in[\frac{1}{2},\infty)$, using $\lim_{\alpha\to 1}$ as $\alpha=1$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]\geq 0$ for an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, and $F^{\beta}_{\alpha}[\mathcal{N}]=0$ iff $\mathcal{N}=\mathcal{T}^{\beta}$. For $\alpha\geq 1$ and $n\in\mathbbm{N}$, $F_{\alpha}^{\beta}[\mathcal{N}^{\otimes n}]=nF^{\beta}_{\alpha}[\mathcal{N}]$ if the outputs of the parallel uses of the channel $\mathcal{N}$ are noninteracting (i.e., the interaction Hamiltonian between the channel outputs is zero). These assertions follow from [30, Theorem 1] and [58, Theorem 4.2].
@@
We note that the expression for the generalized divergence of athermality $\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ was stated in [25, Eqs. (67) & (70)] in relation to its thermal entropy $\mathbf{S}^{\beta}[\mathcal{N}]=-\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$. Albeit in a different context, a quantity termed as the robustness of athermality $R^{\beta}_{T}[\mathcal{N}]$ of a quantum channel, a particular example of the generalized divergence of athermality of the quantum channel, is studied in [59]. The robustness of athermality is directly related to the max-relative entropy of athermality (and max-free energy), $D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]=\ln(1+R^{\beta}_{T})$ (cf. [52, 49, 60, 59]).
@@
The proof of Eq. (9) follows from the fact that $D_{\infty}[\mathcal{N}\|\mathcal{M}]=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\Phi^{\mathcal{M}}_{RA})$ [49] and [25, Eq. (70)],
@@
Details of the proof are in Appendix A.3; the proof for $\alpha=\infty$ follows from [59, Corollary A.4.2], since $F_{\infty}^{\beta}[\mathcal{N}]=\beta^{-1}\ln(1+R^{\beta}_{T}[\mathcal{N}])$.
@@
The proof follows from the definition of the generalized free energy and proof argument similar to [30, Lemma 1]. ∎
@@ IV Dynamical Resource Theory of Athermality
This implies (cf. [64, Proposition 11])
@@
$\qquad=\min_{\Theta}\left\{\frac{1}{2}\|\mathcal{M}-\Theta(\mathcal{N})\|_{\diamond}\penalty 10000\ :\penalty 10000\ \Theta(\mathcal{T}^{\beta}_{1})=\mathcal{T}^{\beta}_{2}\right\}.$ (23)
@@
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (26)
@@
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{1-\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (28)
@@
For any $\varepsilon\in(0,1)$, we have from [65, Theorem 1],
@@
It follows from [65, Theorem 2] that for any athermality distillation scheme which seeks to distill at a rate strictly higher than $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, the probability of successful distillation decays to zero exponentially fast with the number of channel uses of $\mathcal{N}$. That is, $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ also serves the strong converse bound on the asymptotic athermality distillation rate of the channel $\mathcal{N}$, irrespective of whether the strategy used is adaptive or nonadaptive.
@@ IV.1 Remarks on the resource theory of athermality
The dynamical thermodynamic resource theory exhibit some features beyond the the static thermodynamic resource theory. When Hamiltonian $\widehat{H}_{A}$ is non-degenarate, i.e., no two distinct energy eigenstates have same eigenvalues, then a state has the maximum free energy if and only if it is the most excited energy eigenstate (eigenstate with the highest energy eigenvalue). The state with maximum free energy for a quantum system with non-degenerate Hamiltonian is unique. This is characteristically different from the dynamical case where all the unitary channels $\mathcal{U}_{A^{\prime}\to A}$ have maximum free energy. We do have the resource theory of entanglement of states where there are many golden units called maximally entangled states [71], which are all equivalent to each other up to the action of local unitary channels.
@@ V Informational aspects of free energy
Private randomness distillation.— The private randomness capacity of a channel $\mathcal{N}_{A^{\prime}\to A}$ is the maximum rate, in an asymptotic setting of i.i.d. uses of the channel, at which the receiver accessing output $A$ can extract private randomness, against an eavesdropper accessing extension $E$ of an isometric channel extension $\mathcal{V}^{\mathcal{N}}_{A^{\prime}\to AE}$ of the channel N, when the sender sends states through the channel [35]. The private randomness capacity $P[\mathcal{N}]=D[\mathcal{N}\|\mathcal{R}^{\pi}]=\lim_{\beta\to 0^{+}}D[\mathcal{N}\|\mathcal{T}^{\beta}]$. The private randomness capacity of a square quantum channel is twice its asymptotic thermodynamic distillation or formation rate under the action of GPSCs in the limit inverse temperature tends to vanish.
@@
where $Z^{\beta}_{A}=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. For the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]$, where $|A|<\infty$ and $\alpha\in[\frac{1}{2},\infty)$, we have
@@
The min-entropy $S_{\infty}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is associated with the decoupling ability of the channel [38, Theorem 1 & Proposition 1] and the environment-assisted erasure cost $W[A|E]_{\mathcal{N}}$ of the output $A$ of the channel when access to the environment $E$ is provided to the erasure [38, Theorem 3 & Proposition 2]. These operational interpretations of the min-entropy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ directly provides operational interpretations for given information processing tasks as the max-free energy of the channel $F^{\beta}_{\infty}[\mathcal{N}]=\beta^{-1}(\ln|A|-S_{\infty}[\mathcal{N}])$ for $\widehat{H}_{A}=0$.
@@ VI Thermodynamical free energy and work extraction
The function $\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]=-\mathbf{S}^{\beta}[\mathcal{N}]$, where $\mathbf{S}^{\beta}[\mathcal{N}]$ is called the generalized thermal entropy of the channel $\mathcal{N}$ [25]. We see that $\mathcal{F}^{\beta}_{\rm T}[\mathcal{N}]=\beta^{-1}\mathcal{S}^{\beta}[\mathcal{N}]$. The thermodynamic sandwiched Rényi free divergence $F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty]$. The sandwiched Rényi free energy of the absolutely thermal channel reduces to the Helmholtz free energy,
$F^{\beta}_{{\rm T},\alpha}[\widehat{\mathcal{T}}^{\beta}]=-\beta^{-1}\ln Z^{\beta}_{A},$ (55)
@@
The thermodynamic energy $E[\mathcal{N}]$ of a quantum channel $\mathcal{N}$ as
@@ VI.1 Extractable work from a quantum channel
    Reversible, isothermal process: We quasistatically change the Hamiltonian from $\beta^{-1}\ln Z^{\beta}$ back to $\widehat{H}$. Since the state at the initial point of this process is at equilibrium, the state at the final point of the reversible process will be $\gamma_{\beta}$. The work done during this reversible process is given by

### v3

@@ I.1 Main Results
We find that the single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, up to allowed error $\varepsilon\in(0,1)$ under the action of Gibbs preserving superchannels, are $\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\frac{1}{2}D_{H}^{\varepsilon^{2}}[\mathcal{N}\|\mathcal{T}^{\beta}]$, respectively (Theorem 2). We have $F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$, which provide exact operational meaning to the resource-theoretic hypothesis-testing free energy and the max-free energy of a quantum channel. The asymptotic athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, under parallel uses of channels, are equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ (Theorem 3). The resource-theoretic free energy of a channel $\mathcal{N}$, $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, and thus its asymptotic athermality distillation and formation rates are both equal to $\frac{\beta}{2}F^{\beta}[\mathcal{N}]$, providing the exact operational interpretation to the resource-theoretic free energy of a quantum channel. Both asymptotic athermality distillation and formation rates of a quantum channel are the same, implying that the athermality of quantum channels under Gibbs preserving superchannels is (asymptotically) reversible.
@@ II Preliminaries
The generalized channel divergences are monotonically nonincreasing under the action of quantum superchannels [23, Theorem 1]. If the generalized state divergence is nonnegative, then the corresponding channel divergence is also nonnegative. If the generalized state divergence is faithful, i.e., minimum if and only the states are equal, then the corresponding channel divergence is also faithful, i.e., attains minimum if and only if the channels are equal.
@@ III.1 Properties of the dynamical free energy
All the generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ satisfy (A1) and (A2) readily, whenever $\beta^{-1}\mathbf{D}(\rho\|\gamma^{\beta})$ represents some resource-theoretic free energy function $f(\rho)$ for states $\rho$ [14, 44]. The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ remains invariant under the action of Gibbs preserving unitary superchannels. Faithfulness (A3) is satisfied by the generalized free energy whose generalized channel divergence is faithful. For all $\alpha\in[\frac{1}{2},\infty)$, using $\lim_{\alpha\to 1}$ as $\alpha=1$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]\geq 0$ for an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, and $F^{\beta}_{\alpha}[\mathcal{N}]=0$ iff $\mathcal{N}=\mathcal{T}^{\beta}$. For $\alpha\geq 1$ and $n\in\mathbbm{N}$, $F_{\alpha}^{\beta}[\mathcal{N}^{\otimes n}]=nF^{\beta}_{\alpha}[\mathcal{N}]$ if the outputs of the parallel uses of the channel $\mathcal{N}$ are noninteracting (i.e., the interaction Hamiltonian between the channel outputs is zero). These assertions follow from [23, Theorem 1] and [38, Theorem 4.2].
@@
We note that the expression for the generalized divergence of athermality $\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ was stated in [85, Eqs. (67) & (70)] in relation to its thermal entropy $\mathbf{S}^{\beta}[\mathcal{N}]=-\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$. Albeit in a different context, a quantity termed as the robustness of athermality $R^{\beta}_{T}[\mathcal{N}]$ of a quantum channel, a particular example of the generalized divergence of athermality of the quantum channel, is studied in [65]. The robustness of athermality is directly related to the max-relative entropy of athermality (and max-free energy), $D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]=\ln(1+R^{\beta}_{T})$ (cf. [29, 58, 63, 65]).
@@
The proof of Eq. (9) follows from the fact that $D_{\infty}[\mathcal{N}\|\mathcal{M}]=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\Phi^{\mathcal{M}}_{RA})$ [58] and [85, Eq. (70)],
@@
Details of the proof are in Appendix A.3; the proof for $\alpha=\infty$ follows from [65, Corollary A.4.2], since $F_{\infty}^{\beta}[\mathcal{N}]=\beta^{-1}\ln(1+R^{\beta}_{T}[\mathcal{N}])$.
@@
The proof follows from the definition of the generalized free energy and proof argument similar to [23, Lemma 1]. ∎
@@ IV Dynamical Resource Theory of Athermality
This implies (cf. [24, Proposition 11])
@@
$\qquad=\min_{\Theta}\left\{P[\mathcal{M},\Theta(\mathcal{N})]\penalty 10000\ :\penalty 10000\ \Theta(\mathcal{T}^{\beta}_{1})=\mathcal{T}^{\beta}_{2}\right\}.$ (23)
Here $P[\mathcal{M},\mathcal{K}]$ is the purified channel distance between the channels $\mathcal{M}$ and $\mathcal{K}$ given by
$P[\mathcal{M},\mathcal{K}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}P(\mathcal{M}(\psi_{RA^{\prime}}),\mathcal{K}(\psi_{RA^{\prime}}))$ (24)
where $P(\textperiodcentered,\textperiodcentered)$ is the purified distance, and it suffices to take $\psi_{RA^{\prime}}$ to be pure.
@@
$\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ $=\frac{1}{2}D_{H}^{\varepsilon^{2}}[\mathcal{N}\|\mathcal{T}^{\beta}],$ (27)
@@
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
$\mathrm{Cost}^{\sqrt{\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})$ $\leq\mathrm{Dist}^{\sqrt{1-\varepsilon}}(\mathcal{N},\mathcal{T}^{\beta})+\frac{1}{2}\ln\left(\frac{1}{1-\varepsilon}\right).$ (29)
@@
For any $\varepsilon\in(0,1)$, we have from [21, Theorem 1],
@@
It follows from [21, Theorem 2] that for any athermality distillation scheme which seeks to distill at a rate strictly higher than $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, the probability of successful distillation decays to zero exponentially fast with the number of channel uses of $\mathcal{N}$. That is, $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ also serves the strong converse bound on the asymptotic athermality distillation rate of the channel $\mathcal{N}$, irrespective of whether the strategy used is adaptive or nonadaptive.
@@ IV.1 Remarks on the resource theory of athermality
The dynamical thermodynamic resource theory exhibit some features beyond the static thermodynamic resource theory. When Hamiltonian $\widehat{H}_{A}$ is non-degenarate, i.e., no two distinct energy eigenstates have same eigenvalues, then a state has the maximum free energy if and only if it is the most excited energy eigenstate (eigenstate with the highest energy eigenvalue). The state with maximum free energy for a quantum system with non-degenerate Hamiltonian is unique. This is characteristically different from the dynamical case where all the unitary channels $\mathcal{U}_{A^{\prime}\to A}$ have maximum free energy. We do have the resource theory of entanglement of states where there are many golden units called maximally entangled states [51], which are all equivalent to each other up to the action of local unitary channels.
@@ V Informational aspects of free energy
Private randomness distillation.— The private randomness capacity of a channel $\mathcal{N}_{A^{\prime}\to A}$ is the maximum rate, in an asymptotic setting of i.i.d. uses of the channel, at which the receiver accessing output $A$ can extract private randomness, against an eavesdropper accessing extension $E$ of an isometric channel extension $\mathcal{V}^{\mathcal{N}}_{A^{\prime}\to AE}$ of the channel $\mathcal{N}$, when the sender sends states through the channel [94]. The private randomness capacity $P[\mathcal{N}]=D[\mathcal{N}\|\mathcal{R}^{\pi}]=\lim_{\beta\to 0^{+}}D[\mathcal{N}\|\mathcal{T}^{\beta}]$. The private randomness capacity of a square quantum channel is twice its asymptotic thermodynamic distillation or formation rate under the action of GPSCs in the limit inverse temperature tends to vanish.
@@
where $Z^{\beta}_{A}=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. For the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]$, where $|A|<\infty$ and $\alpha\in[\frac{1}{2},\infty)$, we have
@@
The min-entropy $S_{\infty}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is associated with the decoupling ability of the channel [4, Theorem 1 & Proposition 1] and the environment-assisted erasure cost $W[A|E]_{\mathcal{N}}$ of the output $A$ of the channel when access to the environment $E$ is provided to the erasure [4, Theorem 3 & Proposition 2]. These operational interpretations of the min-entropy of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ directly provides operational interpretations for given information processing tasks as the max-free energy of the channel $F^{\beta}_{\infty}[\mathcal{N}]=\beta^{-1}(\ln|A|-S_{\infty}[\mathcal{N}])$ for $\widehat{H}_{A}=0$.
@@ VI Thermodynamical free energy and work extraction
The function $\mathbf{D}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]=-\mathbf{S}^{\beta}[\mathcal{N}]$, where $\mathbf{S}^{\beta}[\mathcal{N}]$ is called the generalized thermal entropy of the channel $\mathcal{N}$ [85]. We see that $\mathcal{F}^{\beta}_{\rm T}[\mathcal{N}]=\beta^{-1}\mathcal{S}^{\beta}[\mathcal{N}]$. The thermodynamic sandwiched Rényi free energy $F^{\beta}_{{\rm T},\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$ for $\alpha\in[\frac{1}{2},1)\cup(1,\infty]$. The sandwiched Rényi free energy of the absolutely thermal channel reduces to the Helmholtz free energy,
$F^{\beta}_{{\rm T},\alpha}[\mathcal{T}^{\beta}]=-\beta^{-1}\ln Z^{\beta}_{A},$ (56)
@@
The thermodynamic energy $E[\mathcal{N}]$ of a quantum channel $\mathcal{N}$ is defined as
@@ VI.1 Extractable work from a quantum channel
    Reversible, isothermal process: We quasistatically change the Hamiltonian from $\beta^{-1}\ln\rho$ back to $\widehat{H}$. Since the state at the initial point of this process is at equilibrium, the state at the final point of the reversible process will be $\gamma_{\beta}$. The work done during this reversible process is given by